OpenStudy (anonymous):

Need help with an equation! Solve the equation x(m+7)+y(m-2)=z(m-9) for m, assuming that x+y-z is not equal to 0. Please show me step by step not just the answer

OpenStudy (asnaseer):

you need to collect the terms involving "m". so first step would be to expand the brackets.

OpenStudy (asnaseer):

can you do that?

OpenStudy (anonymous):

like multiply it out?

OpenStudy (asnaseer):

yes

OpenStudy (anonymous):

Yep gimme a second

OpenStudy (asnaseer):

ok :)

OpenStudy (anonymous):

xm+7x+ym-2y=zm-9z

OpenStudy (asnaseer):

good, now try to get all terms involving "m" to the left-hand-side and all the rest of the terms on the right-hand-side of the equals sign

OpenStudy (anonymous):

xm+ym-zm=-9z-7x+2y Not sure if I did that right

OpenStudy (asnaseer):

looks good to me. :) next - factor the m out on the left-hand-side

OpenStudy (anonymous):

m(x+y-z)=-9z-7x+2y

OpenStudy (asnaseer):

perfect - now divide both sides by (x+y-z)

OpenStudy (anonymous):

m=(-9z-7x+2y)/(x+y-z)

OpenStudy (asnaseer):

thats it! can you see why the question stated: "assuming that x+y-z is not equal to 0" ?

OpenStudy (anonymous):

Hmm, I'm still not sure why it said that

OpenStudy (asnaseer):

think about this: x = something / 0 = ?

OpenStudy (asnaseer):

I meant to say: m = something / 0 = ?

OpenStudy (asnaseer):

if you have an equation like this: 17 * 0 = 16 - 15 - 1 then it does NOT imply that: 17 = (16 - 15 - 1)/0

OpenStudy (asnaseer):

because ANYTHING divided by zero is infinity

OpenStudy (asnaseer):

in algebra you cannot divide by zero - that is not a legal operation to perform

OpenStudy (anonymous):

Ah I think I see what you're saying now

OpenStudy (asnaseer):

great! just one more point - the "standard" way of writing the answer is as follows: m=(-7x+2y-9z)/(x+y-z) i.e try to list the unknowns in alphabetical order

OpenStudy (asnaseer):

what you wrote is still correct - what I suggested is just the "normal" practice in algebra

OpenStudy (asnaseer):

its not a hard and fast rule though

OpenStudy (anonymous):

Oh I didn't know that. =p Thanks so much for the help, I was stuck on that problem for a while. I get lost in all the letters

OpenStudy (asnaseer):

yw :)